# BeSplat – Gaussian Splatting from a Single Blurry Image and Event Stream

### v1

@@ 5.2 Qualitative Evaluations
The qualitative evaluation results, presented in Fig. 3 and Fig. 4, demonstrate the performance of our method on both synthetic and real datasets. Our approach delivers results comparable to BeNeRF while offering significant benefits, including faster training times, real-time rendering capabilities, and reduced GPU memory usage. Specifically, Fig. 3 showcases how prior learning-based methods often struggle to generalize effectively, whereas our method consistently preserves reconstruction quality across diverse scenarios. Additionally, Fig. 4 highlights the robustness of our approach on real noisy datasets, achieving competitive results and maintaining high visual fidelity despite challenging conditions.
@@ 6 Conclusion
We introduce a novel method based on the Gaussian Splatting framework that simultaneously recovers the underlying 3D scene representation and estimates the precise camera motion trajectory from a single blurry image and its associated event stream. Through extensive evaluations on synthetic and real-world datasets, our approach demonstrates a compelling balance of efficiency and quality, achieving accelerated training times, reduced GPU memory usage, and real-time rendering capabilities, all while delivering high-quality reconstructions that validate its practical robustness and versatility. By leveraging the complementary event stream data, our method effectively captures motion details and enhances reconstruction fidelity. Future work could explore extending the framework to handle dynamic scenes, allowing robust reconstruction in environments with non-static objects.
Acknowledgement: We would like to acknowledge partial support from the Institute Of Eminence (IoE) Research Center on VR and Haptics. KM would like to acknowledge support from Qualcomm Faculty Award 2024.

### v2

# BeSplat: Gaussian Splatting from a Single Blurry Image and Event Stream

@@ 5.2 Qualitative Evaluations
The qualitative evaluation results, presented in Fig. 3 and Fig. 4, demonstrate the performance of our method on both synthetic and real datasets.Although BeNeRF, as a NeRF-based approach, achieves slightly better performance due to its detailed neural representation, our method, which leverages Gaussian Splatting, delivers high-quality results with the added benefits of faster training times, real-time rendering capabilities, and reduced GPU memory usage. Specifically, Fig. 3 showcases how our method consistently preserves reconstruction quality across diverse scenarios. Additionally, Fig. 4 highlights the robustness of our approach on real noisy datasets, achieving competitive results and maintaining high visual fidelity despite challenging conditions.
@@ 6 Conclusion
We introduce a novel method based on the Gaussian Splatting framework that simultaneously recovers the underlying 3D scene representation and estimates the precise camera motion trajectory from a single blurry image and its associated event stream. Through extensive evaluations on synthetic and real-world datasets, our approach demonstrates a compelling balance of efficiency and quality, achieving accelerated training times, reduced GPU memory usage, and real-time rendering capabilities, all while delivering high-quality reconstructions that validate its practical robustness and versatility. By leveraging the complementary event stream data, our method effectively captures motion details and enhances reconstruction fidelity. Future work could explore extending the framework to handle dynamic scenes.
Acknowledgement: We would like to acknowledge partial support from the Institute of Eminence (IoE) Research Center on VR and Haptics, together with the IITM Pravartak Technologies Foundation. KM would like to acknowledge support from Qualcomm Faculty Award 2024.